What is the wavelength of a 3128.8 Hz pitch?
1
Expert's answer
2019-02-22T09:52:28-0500

We can find the wavelength of the pitch from the formula:


"\lambda = \dfrac{c}{f},"


here, "\lambda" is the wavelength of the pitch, "c" is the speed of the light and "f" is the frequency of the pitch.

Then, we get:


"\lambda = \dfrac{c}{f} = \dfrac{3 \cdot 10^8 \dfrac{m}{s}}{3128.8 Hz} = 9.58 \cdot 10^4 m."

Answer:

"\lambda = 9.58 \cdot 10^4 m."